OpenStudy (anonymous):

What is the f'(x) of f(X)=sin^3x

OpenStudy (anonymous):

Have you attempted this one yet ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[f(x) = \sin ^{3}(x)\] Did I write this down correctly? If so, try substitution. For example, let u = sin(x) and then take the derivative of u^3

OpenStudy (anonymous):

I got f'(X) is 9x^2cosx^2

OpenStudy (anonymous):

that is 3u^2

OpenStudy (anonymous):

3u^2 du/dx, yes. Now what is du/dx if u = sin x?

OpenStudy (anonymous):

In other words, what is u' if u = sin x

OpenStudy (anonymous):

cosx^2

OpenStudy (anonymous):

What is the derivative of sin(x) ?

OpenStudy (anonymous):

COS

OpenStudy (anonymous):

Great, so du/dx = cos x So we have 3u^2 * cos x Substitute u = sin(x) back in and what do we end up with ?

OpenStudy (anonymous):